# IQ Skew and Imbalance Estimation for Coherent Point-to-Multi-Point Optical Networks

### 31 Jan 2024 v1

@@ I Introduction
Figure 1 shows the schematic diagram of coherent PtMP optical networks based on DSCM with Tx IQ impairments. The hub transmits the DSCM signal with four subcarriers to the leaves, and one leaf receives its corresponding subcarrier rather than symmetrical subcarrier pairs. The above-mentioned schemes for coherent P2P optical networks require the whole spectrum of the transmitted signal to estimate the Tx IQ impairments, which are difficult to apply in DSCM-based coherent PtMP optical networks. This paper proposes the first far-end IQ-impairments estimation for the DSCM-based coherent PtMP optical networks, which can accurately estimate IQ skews and power imbalances of the hub Tx and leaf receiver (Rx) at the leaf. The main contributions of the work are as follows:
@@
The remainder of the paper is organized as follows. In Section II, the TFITs structure is designed to estimate IQ impairments. For DSCM-based coherent PtMP optical networks, the principle of IQ-impairments estimation is introduced. In Section III, we introduce the experimental setups of a DSCM-based coherent PtMP optical network. In Section IV, the experimental results and discussions are demonstrated to verify the performance of the TFITs-based IQ-impairments estimation. Finally, the paper is concluded in Section V.
@@ II-A TFITs Design for IQ-impairments estimation
Specifically, $f_{1}$ and $f_{2}$ can be set to be a quarter of the baud rate $f_{s}/4$ and half of the baud rate $f_{s}/2$, respectively where $f_{s}/4$ denotes the baud rate. Therefore, by combining the Eq. (6) and Eq. (7), the TFITs in the time slot $t_{1}$ can be expressed as
@@
The TFITs can be generated by combining the signals of Eq. (8) and Eq. (9). Fig. 3 shows the specially designed TFITs include four time slots [$t_{1}$, $t_{2}$, $t_{3}$, $t_{4}$] and two tones at frequencies [$f_{1}$, $f_{2}$] for (a) X and (b) Y polarizations. At the $t_{1}$ and $t_{2}$ time slots, the TFITs are carried on the X polarization, while the Y polarization is zero to avoid inter-polarization crosstalk between X and Y polarizations [24, 25, 26]. At the $t_{3}$ and $t_{4}$ time slots, the proposed TFITs are carried on the Y polarization, while the X polarization is zero.
@@ II-B The Model of IQ Impairments on TFITs
where $s_{\text{Rx},f_{s}/2}(t)$ and $js_{\text{Rx},f_{s}/4}(t)$ are denoted as Eq. (13) and Eq. (14) show at the top of this page, respectively. $\phi_{1}(t)$ is $-(2\pi\Delta ft+\phi)$ and $\phi_{2}(t)$ is $-(2\pi\Delta ft+\phi-2\pi f_{\text{c}}\tau_{\text{Tx}})$. After frequency downshift, both the I and Q tributaries of the received signal $s_{\text{Rx}}(t)$ contain the $f_{s}/4$ and $f_{s}/2$ tones, which can be expressed as
@@
where $r_{f_{s}/2,\text{I}}(t)$, $r_{f_{s}/4,\text{I}}(t)$, $r_{f_{s}/2,\text{Q}}(t)$ and $r_{f_{s}/4,\text{Q}}(t)$ are denoted as Eq. (13) and Eq. (14) shown. Therefore, the I and Q tributaries of the signal $s_{\text{Rx}}(t)$ can be expressed as
@@ II-C1 Estimation of Rx IQ Skew
By substituting Eq. (13) and Eq. (14) into Eq. (18), it is inferred that in the time slot $t_{1}$, $r_{f_{s}/2,\text{I}}(t)$ contains only the timing phase offset caused by the out-of-synchronization clock, while $g_{\text{Rx}}r_{f_{s}/2,\text{Q}}(t+\tau_{\text{Rx}})$ contains both the timing phase offset of the clocks and Rx IQ skew $\tau_{\text{Rx}}$. Similarly, there is a difference of Rx IQ skew $\tau_{\text{Rx}}$ between the timing phase offsets of $r_{f_{s}/4,\text{I}}(t)$ and $g_{\text{Rx}}r_{f_{s}/4,\text{Q}}(t+\tau_{\text{Rx}})$. Therefore, the Rx IQ skew can be estimated by calculating the difference in the timing phase offset between the Q and the I tributaries of the tones with the same frequency. The natural high signal-to-noise ratio (SNR) of the frequency tones results in a high estimation accuracy.
Figure 4(a) shows the schematic diagram of Rx IQ-skew estimation by using the I and Q tributaries of the Rx TFITs. By performing sampling and fast Fourier transform (FFT) on the I tributary of the received signal $s_{\text{Rx, I}}(t)$ to $S_{\text{Rx, I}}(k)$ and Q tributary $g_{\text{Rx}}s_{\text{Rx, Q}}(t+\tau_{\text{Rx}})$ to $S_{\text{Rx, Q}}(k)$ at each time slot, the timing phase offsets of I and Q tributaries can be calculated by the Godard phase detector using the $2n$ frequency points around $f_{s}/2$ [30]. Therefore, the Rx IQ skew is estimated by calculating the difference between timing phase offsets of the Q and I tributaries of $f_{s}/2$ tone, which can be expressed as
@@
where the FFT points of $S_{\text{Rx, I}}(k)$ and $S_{\text{Rx, Q}}(k)$ around the $83N/$ 8 correspond to the frequency points around $f_{s}/4$. Finally, the more accurate Rx IQ skew $\tau_{\text{Rx,IQ}}$ can be obtained by averaging $\tau_{\text{Rx,}f_{s}/2}$ and $\tau_{\text{Rx,}f_{s}/4}$.
@@ II-C2 Estimation of Rx IQ Power Imbalance
$C_{f_{s}/4}=2E\left[I_{f_{s}/4}(t+\tau_{\text{Tx}})Q_{f_{s}/4}(t+\tau_{\text{Tx}})\sin 2\phi_{2}(t)\right]$ (27)
@@
Figure 4(b) shows the schematic diagram for Rx IQ power imbalance estimation by using the I and Q tributaries of the Rx TFITs. After FFT, two $f_{s}/2$ tones and two $f_{s}/4$ tones on the I and Q tributaries are extracted, respectively. According to Paswal’s theorem, $g_{\text{Rx,}f_{s}/2}^{2}$ can be calculated by the power ratio between the two $f_{s}/2$ tones extracted from I and Q tributaries. Meanwhile, $g_{\text{Rx,}f_{s}/4}^{2}$ can be calculated by the power ratio between the extracted two $f_{s}/4$ tones from I and Q tributaries. It can improve the accuracy by filtering out the noise at the frequencies outside $f_{s}/2$ and $f_{s}/4$. Finally, the more accurate Rx IQ power imbalance $g^{2}_{\text{Rx,IQ}}$ can be obtained by averaging the $g_{\text{Rx,}f_{s}/2}^{2}$ and $g_{\text{Rx,}f_{s}/4}^{2}$.
@@ II-D1 Estimation of Tx IQ Skew
Figure 5(a) shows the schematic diagram of Tx IQ skew estimation based on the TFITs. By performing sampling and FFT on the Rx signal $s_{\text{Rx},t_{1}}(t)$ in the time slots $t_{1}$ to $S_{\text{Rx},t_{1}}(k)$ and $s_{\text{Rx},t_{2}}(t)$ in the time slots $t_{2}$ to $S_{\text{Rx},t_{2}}(k)$, the timing phase offset of $f_{s}/2$ and $f_{s}/4$ tones are estimated by the Godard phase detector and the modified Godard phase detector, respectively. Thus, the Tx IQ skew can be estimated by subtracting the timing phase offset of the $f_{s}/2$ and $f_{s}/4$ tones in the time slot $t_{1}$ as
@@ III Experimental Setups
Figure 6 shows the experimental setups of an 8Gbaud/SC×$\times$4SCs DSCM-based coherent PtMP optical network to verify the feasibility of the TFITs-based far-end IQ-impairments estimation. At the hub, the bit sequence was mapped to 16 quadrature amplitude modulation (16QAM) symbols. Then a root-raised cosine (RRC) pulse shaping filter with a roll-off factor of 0.1 was used for each subcarrier. After the frequency shift, the four subcarriers were multiplexed to generate a DSCM signal. The specially designed IQ TFITs were inserted before the DSCM signal for estimating the IQ impairments. A pre-compensation of Tx IQ impairments for the DSCM signal was implemented at the hub based on the far-end estimated IQ impairments at the leaf.
@@
To verify the universality of the TFITs-based far-end IQ-impairments estimation, pre-set Tx IQ impairments were added to the digital DSCM signal. After adding pre-set Tx IQ impairments, the digital signal was converted to the analog signal by a digital-to-analog converter (DAC) with a sampling rate of 90GSa/s. Inset (i) in Fig. 6 shows the electrical spectrum of the DSCM signal at the hub. An external cavity laser (ECL) operating at $1550.12$nm with a less than 100kHz linewidth was used as the optical carrier. After being amplified by electrical amplifiers (EAs), the analog signal drove a dual-polarizations IQ modulator (DP I/Q Mod.) to generate the optical signal. The output optical power was approximately $-14$dBm. After being boosted by an Erbium-doped fiber amplifier (EDFA) to approximately 6 dBm, the optical signal was launched into a 100 km standard single-mode fiber (SSMF).
@@
Insets (ii) and (iii) in Fig. 6 are the received spectra of the SC-1 and SC-2 at the leaf, respectively. Without loss of generality, SC-1 and SC-2 were used to estimate the IQ impairments. After frame detection, Rx IQ skew and power imbalance were estimated based on the specially designed TFITs and proposed algorithms. The Rx IQ power imbalance can be compensated by multiplying an estimated coefficient $1/g_{\text{Rx,IQ}}$ in Eq. (21) on the Q tributary. Meanwhile, the Rx IQ skew can be effectively compensated by interpolation based on $\tau_{\text{Rx,IQ}}$. After compensating Rx IQ impairments, the Tx IQ skew $\tau_{\text{Tx,IQ}}$ and power imbalance $g_{\text{Tx,IQ}}$ were estimated by the specially designed TFITs and proposed algorithms and then transmitted to the hub using the uplink. After the IQ impairments estimation, the DSCM signal was recovered by DSP as demonstrated in Refs. [33, 34, 35].
@@ IV Experimental Results and Discussions
Figures 8(a) and (b) show the estimated Tx IQ skews versus pre-set Tx IQ skews and the absolute errors between the estimated Tx IQ skews and pre-set Tx IQ skews for the X and Y polarizations, respectively. When only the Tx IQ skews were set to the values from $-15$ps to $15$ps on X and Y polarizations, the TFITs-based far-end Tx IQ-skew estimation can achieve the absolute error within ±$\pm 0.5$ps. Figs. 8(c) and (d) depict the estimated Tx IQ power imbalances versus pre-set Tx IQ power imbalances and the absolute errors between the estimated Tx IQ power imbalances and pre-set Tx IQ power imbalances for the X and Y polarizations, respectively. When the Tx IQ power imbalances were set to the values from $-3$dB to $3$dB, the TFITs-based far-end Tx IQ-power-imbalance estimation can achieve an absolute error within ±$\pm 0.2$dB.
@@
Figure 9(c) shows the estimated Tx IQ skew versus pre-set Tx IQ skew and the absolute error between the estimated Tx IQ skew and pre-set Tx IQ skew when $1$dB Tx IQ power imbalance, $-5$ps Rx IQ skew, and $-1$dB Rx IQ power imbalance coexist. The Tx IQ skew from $-15$ps to $15$ps can be estimated with the absolute error within ±$\pm 0.5$ps by TFITs-based far-end Tx IQ-skew estimation. Fig. 9(d) shows the estimated Tx IQ power imbalance versus pre-set IQ power imbalance and the absolute error between estimated IQ power imbalance and pre-set IQ power imbalance when $5$ps Tx IQ skew, $-5$ps Rx IQ skew, and $-1$dB Rx IQ power imbalance coexist. The Tx IQ power imbalance from $-3$dB to $3$dB can be estimated with the absolute error within ±$\pm 0.2$dB by TFITs-based far-end Tx IQ-power-imbalance estimation. In conclusion, TFITs-based IQ-impairments estimation can accurately estimate Rx and Tx IQ impairments at the far-end leaf.
@@
Figures 10(c) and (d) depict the BERs of the SC-1 and SC-2 under the Tx IQ skews and Tx IQ power imbalances, respectively. Insets of Figs. 10(c) and (d) show the constellation diagrams under $15$ps Tx IQ skew or $3$dB Tx IQ power imbalance of (i) SC-1 without compensation, (ii) SC-1 with compensation, (iii) SC-2 without compensation, and (iv) SC-2 with compensation, respectively. Since the image crosstalk caused by Tx IQ skew increases with the frequency increase, higher-frequency SC-1 suffers from more distortions than lower-frequency SC-2. Thus, the BER performance of SC-1 is more sensitive to Tx IQ skew than that of SC-2. Since the image crosstalk caused by the Tx IQ power imbalance is the same on all the frequencies, the higher-frequency SC-1 suffers from almost the same distortions as the lower-frequency SC-2, leading to the same trend of BER performance degradation. After compensation for Tx IQ impairments, the BER penalty on the two subcarriers can be almost eliminated, and the constellations become much clearer. In conclusion, the proposed IQ-impairments estimation and compensation can effectively improve the BER performance for the DSCM-based coherent PtMP optical networks.

### 11 Apr 2024 v2

@@ I Introduction
Fig. 1 shows the schematic diagram of coherent PtMP optical networks based on DSCM with Tx IQ impairments. The hub transmits the DSCM signal with four subcarriers to the leaves, and one leaf receives its corresponding subcarrier rather than symmetrical subcarrier pairs. The above-mentioned schemes for coherent P2P optical networks require the whole spectrum of the transmitted signal to estimate the Tx IQ impairments, which are difficult to apply in DSCM-based coherent PtMP optical networks. This paper proposes the first far-end IQ-impairments estimation for the DSCM-based coherent PtMP optical networks, which can accurately estimate IQ skews and power imbalances of the hub Tx and leaf receiver (Rx) at the leaf. The main contributions of the work are as follows:
@@
The remainder of the paper is organized as follows. In Section II, the TFITs structure is designed to estimate IQ impairments. The principle of IQ-impairments estimation is introduced for DSCM-based coherent PtMP optical networks. In Section III, we introduce the experimental setups of a DSCM-based coherent PtMP optical network. In Section IV, the experimental results and discussions are demonstrated to verify the performance of the TFITs-based IQ-impairments estimation. Finally, the paper is concluded in Section V.
@@ II-A TFITs Design for IQ-impairments estimation
The $f_{1}$ and $f_{2}$ can be set to any frequencies within $f_{s}$ where $f_{s}$ denotes the baud rate. Specifically, $f_{1}$ and $f_{2}$ are set to a quarter of the baud rate (i.e. $f_{s}/4$) and half of the baud rate (i.e. $f_{s}/2$), respectively. Thus, by combining the Eq. (6) and Eq. (7), the TFITs in the time slot $t_{1}$ can be expressed as
@@
The TFITs can be generated by combining the signals of Eqs. (8) and (9). Fig. 3 shows the specially designed TFITs include four time slots [$t_{1}$, $t_{2}$, $t_{3}$, $t_{4}$] and two tones at frequencies [$f_{1}$, $f_{2}$] for (a) X and (b) Y polarizations. At the time slots of $t_{1}$ and $t_{2}$, the TFITs are carried on the X polarization, while the Y polarization is zero to avoid inter-polarization crosstalk between X and Y polarizations [24, 25, 26]. At the time slots of $t_{3}$ and $t_{4}$, the proposed TFITs are carried on the Y polarization, while the X polarization is zero.
@@ II-B The Model of IQ Impairments on TFITs
where $s_{\text{Rx},f_{s}/2}(t)$ and $js_{\text{Rx},f_{s}/4}(t)$ are denoted as Eqs. (13) and (14) show at the top of this page, respectively. $\phi_{1}(t)$ is $-(2\pi\Delta ft+\phi)$ and $\phi_{2}(t)$ is $-(2\pi\Delta ft+\phi-2\pi f_{\text{c}}\tau_{\text{Tx}})$. After frequency downshift, both the I and Q tributaries of the received signal $s_{\text{Rx}}(t)$ contain the $f_{s}/4$ and $f_{s}/2$ tones, which can be expressed as
@@
where $r_{f_{s}/2,\text{I}}(t)$, $r_{f_{s}/4,\text{I}}(t)$, $r_{f_{s}/2,\text{Q}}(t)$ and $r_{f_{s}/4,\text{Q}}(t)$ are denoted as Eqs. (13) and (14) show. Therefore, the I and Q tributaries of the signal $s_{\text{Rx}}(t)$ can be expressed as
@@ II-C1 Estimation of Rx IQ Skew
By substituting Eqs. (13) and (14) into Eq. (18), it is inferred that in the time slot $t_{1}$, $r_{f_{s}/2,\text{I}}(t)$ contains only the timing phase offset caused by the out-of-synchronization clock, while $g_{\text{Rx}}r_{f_{s}/2,\text{Q}}(t+\tau_{\text{Rx}})$ contains both the timing phase offset of the clocks and Rx IQ skew $\tau_{\text{Rx}}$. Similarly, there is a difference of Rx IQ skew $\tau_{\text{Rx}}$ between the timing phase offsets of $r_{f_{s}/4,\text{I}}(t)$ and $g_{\text{Rx}}r_{f_{s}/4,\text{Q}}(t+\tau_{\text{Rx}})$. Therefore, the Rx IQ skew can be estimated by calculating the difference in the timing phase offset between the Q and the I tributaries of the tones with the same frequency.
Fig. 4(a) shows the schematic diagram of Rx IQ-skew estimation by using the I and Q tributaries of the Rx TFITs. By performing sampling and fast Fourier transform (FFT) on the I tributary of the received signal $s_{\text{Rx, I}}(t)$ to $S_{\text{Rx, I}}(k)$ and Q tributary $g_{\text{Rx}}s_{\text{Rx, Q}}(t+\tau_{\text{Rx}})$ to $S_{\text{Rx, Q}}(k)$ at each time slot, the timing phase offsets of I and Q tributaries can be calculated by the Godard phase detector using the $2n$ frequency points around $f_{s}/2$ [30]. Therefore, the Rx IQ skew is estimated by calculating the difference between timing phase offsets of the Q and I tributaries of $f_{s}/2$ tone, which can be expressed as
@@
where the FFT points of $S_{\text{Rx, I}}(k)$ and $S_{\text{Rx, Q}}(k)$ around the $83N/$ 8 correspond to the frequency points around $f_{s}/4$. It is worth noting that the calculations for timing phase offset should be modified based on the frequencies $f_{1}$ and $f_{2}$. Finally, the more accurate Rx IQ skew $\tau_{\text{Rx,IQ}}$ can be obtained by averaging $\tau_{\text{Rx,}f_{s}/2}$ and $\tau_{\text{Rx,}f_{s}/4}$.
@@ II-C2 Estimation of Rx IQ Power Imbalance
$C_{f_{s}/4}=2E\left[I_{f_{s}/4}(t+\tau_{\text{Tx}})Q_{f_{s}/4}(t+\tau_{\text{Tx}})\sin 2\phi_{2}(t)\right].$ (27)
@@
Fig. 4(b) shows the schematic diagram for Rx IQ power imbalance estimation by using the I and Q tributaries of the Rx TFITs. After FFT, two $f_{s}/2$ tones and two $f_{s}/4$ tones on the I and Q tributaries are extracted, respectively. According to Paswal’s theorem, $g_{\text{Rx,}f_{s}/2}^{2}$ can be calculated by the power ratio between the two $f_{s}/2$ tones extracted from I and Q tributaries. Meanwhile, $g_{\text{Rx,}f_{s}/4}^{2}$ can be calculated by the power ratio between the extracted two $f_{s}/4$ tones from I and Q tributaries. It can improve the accuracy by filtering out the noise at the frequencies outside $f_{s}/2$ and $f_{s}/4$. Finally, the more accurate Rx IQ power imbalance $g^{2}_{\text{Rx,IQ}}$ can be obtained by averaging the $g_{\text{Rx,}f_{s}/2}^{2}$ and $g_{\text{Rx,}f_{s}/4}^{2}$.
@@ II-D1 Estimation of Tx IQ Skew
Fig. 5(a) shows the schematic diagram of Tx IQ skew estimation based on the TFITs. By performing sampling and FFT on the Rx signal $s_{\text{Rx},t_{1}}(t)$ in the time slots $t_{1}$ to $S_{\text{Rx},t_{1}}(k)$ and $s_{\text{Rx},t_{2}}(t)$ in the time slots $t_{2}$ to $S_{\text{Rx},t_{2}}(k)$, the timing phase offset of $f_{s}/2$ and $f_{s}/4$ tones are estimated by the Godard phase detector and the modified Godard phase detector, respectively. Thus, the Tx IQ skew can be estimated by subtracting the timing phase offset of the $f_{s}/2$ and $f_{s}/4$ tones in the time slot $t_{1}$ as
@@ III Experimental Setups
Fig. 6 shows experimental setups of an 8Gbaud/SC×$\times$4SCs DSCM-based coherent PtMP optical network to verify the feasibility of the TFITs-based far-end IQ-impairments estimation. The bit sequence was mapped to 16 quadrature amplitude modulation (16QAM) symbols at the hub. Then, each subcarrier used a root-raised cosine (RRC) pulse-shaping filter with a roll-off factor of 0.1. After the frequency upshift, the four subcarriers were multiplexed to generate a DSCM signal. The specially designed TFITs were inserted before the DSCM signal for estimating the IQ impairments at the expense of spectral efficiency. Owing to the slowly changing IQ impairments, we can transmit the TFITs at intervals such as every online registration, not every data frame. In our experiment, one block of TFITs in Fig. 3 contained 2048 symbols. We averaged the IQ impairments estimated by 3 blocks of TFITs to achieve higher estimation accuracy. A pre-compensation of Tx IQ impairments for the DSCM signal was implemented at the hub based on the far-end estimated IQ impairments at the leaf. The Tx IQ power imbalance can be compensated by multiplying an estimated coefficient $1/g_{\text{Tx,IQ}}$ on the Q tributary. Meanwhile, the Tx IQ skew can be effectively compensated by interpolation based on $\tau_{\text{Tx,IQ}}$.
@@
To verify the universality of the TFITs-based far-end IQ-impairments estimation, pre-set Tx IQ impairments were added to the digital DSCM signal. After adding pre-set Tx IQ impairments, a digital-to-analog converter (DAC) converted the digital signal to an analog signal with a sampling rate of 90GSa/s. Inset (i) in Fig. 6 shows the electrical spectrum of the DSCM signal at the hub. An external cavity laser (ECL) operating at $1550.12$nm with a less than 100kHz linewidth was used as the optical carrier. After being amplified by electrical amplifiers (EAs), the analog signal drove a dual-polarizations IQ modulator (DP I/Q Mod.) to generate the optical signal. The output optical power was approximately $-14$dBm. After being boosted by an Erbium-doped fiber amplifier (EDFA) to approximately 6 dBm, the optical signal was launched into a 100 km standard single-mode fiber (SSMF).
@@
Insets (ii) and (iii) in Fig. 6 are the received spectra of the SC-1 and SC-2 at the leaf, respectively. Without loss of generality, SC-1 and SC-2 were used to estimate the IQ impairments. After frame detection, Rx IQ skew and power imbalance were estimated based on the specially designed TFITs and proposed algorithms. The Rx IQ power imbalance can be compensated by multiplying an estimated coefficient $1/g_{\text{Rx,IQ}}$ in Eq. (21) on the Q tributary. Meanwhile, the Rx IQ skew can be effectively compensated by interpolation based on $\tau_{\text{Rx,IQ}}$. After compensating Rx IQ impairments, the Tx IQ skew $\tau_{\text{Tx,IQ}}$ and power imbalance $g_{\text{Tx,IQ}}$ were estimated by the specially designed TFITs and proposed algorithms and then transmitted to the hub using the uplink. The estimated Tx IQ skews and power imbalances can be averaged at the hub to achieve a more accurate IQ skew and power imbalance. After the IQ impairments estimation, the DSCM signal was recovered by DSP as demonstrated in Refs. [33, 34, 35].
@@ IV Experimental Results and Discussions
Figs. 8(a) and (b) show the estimated Tx IQ skews versus pre-set Tx IQ skews and the absolute errors between the estimated Tx IQ skews and pre-set Tx IQ skews for the X and Y polarizations, respectively. When only the Tx IQ skews were set to the values from $-15$ps to $15$ps on X and Y polarizations, the TFITs-based far-end Tx IQ-skew estimation can achieve the absolute error within ±$\pm 0.5$ps. Figs. 8(c) and (d) depict the estimated Tx IQ power imbalances versus pre-set Tx IQ power imbalances and the absolute errors between the estimated Tx IQ power imbalances and pre-set Tx IQ power imbalances for the X and Y polarizations, respectively. When the Tx IQ power imbalances were set to the values from $-3$dB to $3$dB, the TFITs-based far-end Tx IQ-power-imbalance estimation can achieve an absolute error within ±$\pm 0.2$dB.
@@
Fig. 9(c) shows the estimated Tx IQ skew versus pre-set Tx IQ skew and the absolute error between the estimated Tx IQ skew and pre-set Tx IQ skew when $1$dB Tx IQ power imbalance, $-5$ps Rx IQ skew, and $-1$dB Rx IQ power imbalance coexist. The Tx IQ skew from $-15$ps to $15$ps can be estimated with the absolute error within ±$\pm 0.5$ps by TFITs-based far-end Tx IQ-skew estimation. Fig. 9(d) shows the estimated Tx IQ power imbalance versus pre-set IQ power imbalance and the absolute error between estimated IQ power imbalance and pre-set IQ power imbalance when $5$ps Tx IQ skew, $-5$ps Rx IQ skew, and $-1$dB Rx IQ power imbalance coexist. The Tx IQ power imbalance from $-3$dB to $3$dB can be estimated with the absolute error within ±$\pm 0.2$dB by TFITs-based far-end Tx IQ-power-imbalance estimation. In conclusion, TFITs-based IQ-impairments estimation can accurately estimate Rx and Tx IQ impairments at the far-end leaf.
@@
Figs. 10(c) and (d) depict the BERs of the SC-1 and SC-2 under the Tx IQ skews and Tx IQ power imbalances, respectively. Insets of Figs. 10(c) and (d) show the constellation diagrams under $15$ps Tx IQ skew or $3$dB Tx IQ power imbalance of (i) SC-1 without compensation, (ii) SC-1 with compensation, (iii) SC-2 without compensation, and (iv) SC-2 with compensation, respectively. Since the image crosstalk caused by Tx IQ skew increases with the frequency increase, higher-frequency SC-1 suffers from more distortions than lower-frequency SC-2. Thus, the BER performance of SC-1 is more sensitive to Tx IQ skew than that of SC-2. Since the image crosstalk caused by the Tx IQ power imbalance is the same on all the frequencies, the higher-frequency SC-1 suffers from almost the same distortions as the lower-frequency SC-2, leading to the same trend of BER performance degradation. After compensation for Tx IQ impairments, the BER penalty on the two subcarriers can be almost eliminated, and the constellations become much clearer. In conclusion, the proposed IQ-impairments estimation and compensation can effectively improve the BER performance for the DSCM-based coherent PtMP optical networks.